OpenStudy (anonymous):

What are the intercepts of the equation? Graph the equation. 5x + 3y = 15

OpenStudy (anonymous):

You do the y = this

OpenStudy (anonymous):

3y = 15 - 5x y = 5 - 5/3x

OpenStudy (anonymous):

let x=0, solve for y.... this will give you the y-intercept. let y=0, solve for x... this will give you the x-intercept. draw a line through these two points to show the graph...

OpenStudy (anonymous):

Yes what dpalnc said. Then you just plug into the equation

OpenStudy (anonymous):

so the x intercept is 5,0 and y is 0,3

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

ah thank you so much!